Jose is 3 times as old as wally.Five years from how the sum of their ages is 42.How old are they
Answer by ikleyn(52803)

.
Let x = Wally's age;  then the Joses's age is 3x.


From the condition, you have this equation


    (x+5) + (3x+5) = 42,


    4x = 32

     x = 32/4 = 8.


ANSWER.  Wally is 8 years old.  Jose is 3*8 = 24 years old.

Solved.